OpenStudy (anonymous):

how will I graph ''She is 11 meters into the race after four seconds and is only 13 meters in after 12 seconds''?

OpenStudy (♪chibiterasu):



OpenStudy (♪chibiterasu):

Time will be represented on the x-axis while progression is represented on the y-axis.

OpenStudy (anonymous):

I know the x and y axis thing but I am trying to find the plotting and starting points

OpenStudy (♪chibiterasu):

(x, y) (time, meters)

OpenStudy (anonymous):

i know the x and y but the equation what will it be

OpenStudy (♪chibiterasu):

Well first we find the slope.

OpenStudy (texaschic101):

use your 2 sets of points (4,11) and (12,13)...and find the slope

OpenStudy (♪chibiterasu):

Slope Formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

so what about y=mx+b

OpenStudy (♪chibiterasu):

We can form the equation in slope-intercept form: y = mx + b m = slope b = y-intercept The y-intercept is wher the line crosses the y-axis. Plug that in for b and plug in the slope for m and that is your equation.

OpenStudy (texaschic101):

(4,11)...and (12,13) slope formula : (y2 - y1) / (x2 - x1) (4,11)....x1 = 4 and y1 = 11 (12,13)...x2 = 12 and y2 = 13 now sub that into the slope formula and find the slope

OpenStudy (♪chibiterasu):

And then the y-intercept is just the point where the line connects with the y-axis. That number will be 'b'.

OpenStudy (texaschic101):

so what did you get for the slope ?

OpenStudy (anonymous):

2/8

OpenStudy (texaschic101):

and that reduces to what ?

OpenStudy (anonymous):

1/4

OpenStudy (♪chibiterasu):

And where does the line cross the y-axis?

OpenStudy (texaschic101):

ok...your slope(m) = 1/4 now use y = mx + b slope(m) = 1/4 use either of your points....(4,11)...x = 4 and y = 11 now sub 1/4 in for m and 4 for x and 11 for y into y = mx + b and show me what you get.

OpenStudy (texaschic101):

your basically solving for b, your y intercept

OpenStudy (anonymous):

y=1/4x+b or 11=1/4(4)+b